OpenStudy (anonymous):

Solve the following system of equations by substitution. 3y = 9 and 9y + 2x = -12

OpenStudy (anonymous):

3y = 9 Therefore y = 3 Substituting y = 3 in 9y + 2x = -12, we get (9*3) + 2x = -12 27 + 2x = -12 2x = -12-27 x = -39/2